11. What compound will the following ions form in 5 M NH OH? a. Zn2+ b. Cr3* c. Ni2+ d. Fe*please help me with a lab question about what compounds are formed
Answer #1

Ans Zn2+ + NH4OH = Zn(OH)4 2-

Cr3+ + NH4OH = Cr (OH)3

Ni2+ + NH4OH = Ni(OH)2

Fe3+ + NH4OH = Fe(OH)3